OpenStudy (mackenzic):

will medal the triangles are similar. the area of the larger triangle is 1200cm2 what is the area of the small triangle? the small triangle is 16 cm and the larger triangle is 64cm

OpenStudy (welshfella):

the ratio of the areas = a^2 / b^2 where a and b are the lengths of corresponding sides.

OpenStudy (mackenzic):

im sorry I am still confused on what I have to do

OpenStudy (welshfella):

64^2 / 16^2 is the ratio of the areas ( as i said above - the square of their lengths)

OpenStudy (welshfella):

we are given that the area of large triangle is 1200 cm^2 and if we call the area of the smaller one x we can set up the following proportion 64^2 / 16^2 = 1200 / x solve this for x.

OpenStudy (welshfella):

can you do that?

OpenStudy (mackenzic):

I got 75

OpenStudy (mackenzic):

thank you i got them right :)